The length of a rectangle is 7 less than twice its length. It's area is 30 square centimeters. Find the dimensions of the rectangle.
Answer by stanbon(75887)

The width of a rectangle is 7 less than twice its length.
length = L
width = 2L-7
----------------------------
It's area is 30 square centimeters.
Area = (length)(width) = L(2L-7)
---------------------------------------
Find the dimensions of the rectangle.
Equation:
L(2L-7) = 30
2L^2 - 7L - 30 = 0
-----
2L^2 - 12L+5L - 30 = 0
2L(L-6)+5(L-6) = 0
(L-6)(2L+5) = 0
-----
Positive solution::
Length = 6 cm
Width = 2L-7 = 5 cm
-----------------------
Cheers,
Stan H.
----------------
----------